# DarkDNS: Revisiting the Value of Rapid Zone Update

### v1

@@ 4.2. Transient Domains
While early detection of newly registered domains represents a significant advance detecting possible malicious behavior, those domains will generally show up in the zone snapshots in the days following their registration. However, a small percentage (1%) of newly registered domains never appeared in a zone snapshot: those deleted within the first 24 hours of their existence, which we call transient domains. We identified these names by excluding from our list of newly registered domains those that appeared in our zone collection during the 3 month observation window (+/- 3 days to account for late zone snapshots publishing).
TLD Category Dec Nov Jan Total
com 10 5971059710\,59710 597 $9363$ 21 2322123221\,23221 232 41 1924119241\,19241 192
online $2369$ $1800$ $1990$ $6159$
site $1381$ $1578$ $890$ $3849$
net $866$ $702$ $1544$ $3112$
org $602$ $595$ $1176$ $2373$
shop $497$ $688$ $507$ $1692$
xyz $316$ $321$ $624$ $1261$
store $414$ $422$ $377$ $1213$
top $161$ $213$ $276$ $650$
fun $175$ $185$ $160$ $520$
Others $1958$ $1609$ $2454$ $6021$
Total 19 3361933619\,33619 336 17 4761747617\,47617 476 31 2303123031\,23031 230 68 0426804268\,04268 042
@@
Since domains with non-responding RDAP may not be transient (cause iii), we filtered them from subsequent analysis. Using the RDAP-reported registration timestamp, we also filtered domains misclassified as newly registered. This filtering yielded 42358 confirmed transient domains over the 3-month period.
@@ 4.2.2. Registrar and Hosting Landscape of Transient Domains
Table 3. Transient Domains Registrars Distribution.
The collected RDAP data revealed that the market leader GoDaddy topped the list of registrars holding transient domains, with 19% of the total such domains (Table 3). The registrar landscape is dominated by large registrars, with smaller registrars making up the long tail of the distribution (see Others in Table 3). These results suggest that transient domains are a widespread phenomenon across registrars.
We used our active resolution measurements to analyze the DNS hosting and web hosting infrastructure of transient domains. Half of these transient domains were using Cloudflare as DNS provider (i.e., for their authoritative nameservers) and $\approx$35of them used Cloudflare as a CDN provider. We identified Hostinger’s parking DNS nameservers as the second most prominent DNS provider, as accounting for $\approx$8of the transient domains, followed by a long tail of popular DNS hosting services including NS1, Squarespace, and GoDaddy. These findings suggest that those domains are generally hosted on similar infrastructure of long-lived domains. In contrast, the large presence of parked domains could indicate domains that have already been removed or caught before misuse.
@@ 4.3. Reasons for Early Removals
In private conversations with two top registrars in Table 3, they confirmed that transient domains are mostly likely malicious. Reasons for early removal include abuse, account suspensions, or credit card fraud. To investigate indicators of maliciousness of transient domain names, we analyzed how often blocklists manage to flag them. We considered both early-removed newly registered domain names and 42501 transient domains (Table 3). With early-removed, we refer to newly registered domains that were removed before the end of our analysis period. Unlike transient domains, early-removed domains appear in the zone file snapshots but still have a shorter lifespan than the 1-year typical duration of a registered domain due to the likely forced removal by the registrar. We observed 555491 domains that were deleted before 1 February 2024 (10% of newly registered domains we detected).
@@
Of the 42501 transient domains, at least one blocklist flagged 4.9% (2123) as malicious. Of these, 5% (105) were flagged on their registration date, 1% (12) were already on the blocklists before their registration date, and 94% (2006) appeared after domain deletion.
@@
Blocklists identify and label as malicious a tiny percentage of domains before their registration dates, which may indicate a re-registration. Additionally, a significant percentage (94%) of transient malicious domains appeared on the blocklists post-deletion, indicating that blocklists do not promptly or in some cases ever detect short-lived malicious domains, consistent with other studies (Bouwman et al., 2022).
@@ 4.4. Visibility Gap in Domain Detection
At first, we compared newly registered domains, finding that the SIE NOD feed detects in absolute numbers roughly 5% more domains than our methododology. However, the overlap between the two data sources is $\approx$60%, indicating that both methodologies provide an additional unique subset of domains. Looking at transient domains identified by the two sources, we find the overlap drops further. A total of 855 transient domains were identified by either one of the two sources. Of these, only 33% were detected by both, with the SIE NOD feed identifying 10% more transient domains than our methodology. While the SIE NOD detects a larger number, the clear disjoint nature of the intersecting set of domains suggest the need to combine them to bridge, although not completely, the visibility gap.
To validate our visibility gap against a source of ground truth, we use the perspective of a mid-sized ccTLD registry on transient domains. From November 2023 until January 2024, the registry observed 714 domain names that were deleted in less than 24 hours in their registration system. Of those domains, 334 were registered and deleted such that they were never captured in zone file snapshots. Applying our methodology to this ccTLD, we found only 99 transient domains, or 29.6% of the 334 ccTLD-identified transient domain names. [^5] We were unable to compare the SIE NOD feed to the registry data due to data availability issues. However, based on our previous results we argue that combining the two data sources would still leave a large visibility gap. This result shows that while we can catch a fraction of these transient domains, researchers still have a huge blindspot regarding intra-day events.
@@ 5. Discussion and Conclusion
In the meantime, transient domains, in which malicious activity tends to dominate, have largely been invisible to the research community. At the same time, research has established that the existing DNS-related threat intelligence community, e.g., blocklist companies, are less effective against new and emerging threats (Bouwman et al., 2022). As such, there exists a need for new approaches to address this issue.
@@
Verisign long ago supported rapid zone updates, enabling updates to the .com and .net zone every 5 minutes. The data included domain names, nameservers, IP address additions, deletions and modifications, to “promote security and stability by providing a useful tool to online security companies, ISPs, search engines, financial services companies, and other stakeholders.” (ICANN, 2007) [^6] Appendix A excerpts Verisign’s registry service application.. This service illuminated the activity of bad actors, including phishing, fraud and identity theft (Hao et al., 2013). Concerns about abuse of the data by spammers led to Verisign’s termination of this service. In particular, attackers exploited a brief time window where the zone updates were publicly available, but before domain owners had set up protection of their domain.
@@
Given the inherent advantages that attackers have over defenders, the ineffectiveness of existing uncoordinated countermeasures, and the limited obligations (or ability) of registrars to mitigate harm, [^7] ICANN’s recently revised contract language does not attribute any duty of care to registrars or registries, and does not address vulnerabilities such as exact-match brand registration and suspicious bulk registration (Dave Piscatello and Colin Strutt, 2023). we see a need to expand transparency and accountability mechanisms.
Registrars and registries who want to establish themselves as serious about security could resurrect the capability to subscribe to rapid zone updates along with a framework to safeguard against abuses, and to learn from history how to mitigate the risk of abuse. While risks of abuse of access to non-public DNS data has created tension and controversy for decades, the CZDS program is testament to the ability to manage these risks. More recently, ICANN has coordinated a model for requesting access to other non-public registration data in a more consistent and standardized format (ICANN, 2023). Such access to rapid zone updates would enable a source of labeled data that would allow application of machine learning techniques not only to security research on systemic harms, but also to anti-abuse efforts of registries, registrars, and law enforcement agencies.
@@ Appendix A Ethics
Our ethics review board does not require a specific ethics approval for internet measurements conducted according to community best practices. In measuring the DNS infrastructure of newly registered domains, we followed these best practices by aiming to strike a balance between coverage and slower probing rate. As a result,we decided to scan those domains every 10 minutes. Furthermore, we ensured that the address space of our measurement infrastructure had pointers to direct interested users to a page explaining the project’s goals and providing a contact for opting out from the measurements.
@@ Appendix B Appendix: Verisign’s request to provide Rapid Zone updates
The following is a quotation from Verisign’s application to ICANN to provide rapid zone updates as a service citevrsn-rzu-2007:
@@
  • The service would shed light on the activity of those engaged in domain tasting and expose bad actors. “This [service] elevates bad actors into the light of day.” “All it does is ‘out’ people who are probably bad guys.”
@@
External counsel and ICANN staff agreed that there are no apparent competition issues that require that this matter be forwarded to relevant competition authorities at this time. citevrsn-rzu-2007

### v2

@@ 4.2. Transient Domains
While early detection of newly registered domains represents a significant advance detecting possible malicious behavior, those domains will generally show up in the zone snapshots in the days following their registration. However, a small percentage (1%) of newly registered domains never appeared in a zone snapshot due to their short lifespan falling between two snapshots. We call these domains transient domains. We identified these names by excluding from our list of newly registered domains those that appeared in our zone collection during the 3 month observation window (+/- 3 days to account for late zone snapshots publishing).
@@
TLD Category Nov Dec Jan Total
com $9363$ 10 5971059710\,59710 597 21 2322123221\,23221 232 41 1924119241\,19241 192
online $1800$ $2369$ $1990$ $6159$
site $1578$ $1381$ $890$ $3849$
net $702$ $866$ $1544$ $3112$
org $595$ $602$ $1176$ $2373$
shop $688$ $497$ $507$ $1692$
xyz $321$ $316$ $624$ $1261$
store $422$ $414$ $377$ $1213$
top $213$ $161$ $276$ $650$
fun $185$ $175$ $160$ $520$
Others $1609$ $1958$ $2454$ $6021$
Total 17 4761747617\,47617 476 19 3361933619\,33619 336 31 2303123031\,23031 230 68 0426804268\,04268 042
@@
Since domains with non-responding RDAP may not be transient (cause iii), we filtered them from subsequent analysis. Using the RDAP-reported registration timestamp, we also filtered domains misclassified as newly registered. This filtering yielded 42358 confirmed transient domains over the 3-month period. These transient domains represent only a portion of all the transient domains across the zones we analyzed. Because we do not have access to ground truth (i.e., registries view), we cannot establish the total number of transient domains or the extent of our methodology’s coverage. In subsection 4.4, we will compare our detection with a passive DNS source and the ground truth from a ccTLD.
@@ 4.2.2. Registrar and Hosting Landscape of Transient Domains
The collected RDAP data revealed that market leader GoDaddy topped the list of registrars holding transient domains, with 19% of such domains (Table 3). The registrar landscape is dominated by large registrars, with smaller registrars making up the long tail of the distribution (see Others in Table 3). These results suggest that transient domains are a widespread phenomenon across registrars.
Table 4. Top 5 DNS Hosting Domains (NS records) of Transient Domains
Name NS Record SLD Domains %
Cloudflare cloudflare.com 20981 49.5%
Hostinger dns-parking.com 3682 8.7%
NS1 nsone.net 2938 6,9%
Squarespace squarespacedns.com 2908 6.9%
GoDaddy domaincontrol.com 2315 5.5%
Others - 9534 22.5%
Total - 42358 -
Table 5. Top 5 Web Hosting (A records) of Transient Domains
Name ASN Domains %
Cloudflare 13335 15322 36.2%
Hostinger 47583 5930 14.0%
Amazon 16509 3198 7.6%
Squarespace 53831 2257 5.3%
Namecheap 22612 1650 3.9%
Others - 14001 33.1%
Total - 42358 -
We used our active resolution measurements to analyze the DNS hosting (Table 4) and web hosting (Table 5) infrastructure of transient domains. For DNS hosting, we investigated the most popular nameservers SLDs, while for Web hosting we looked at the ASNs of the A records of the domains. Half of these transient domains were using Cloudflare as DNS provider (i.e., for their authoritative nameservers) and $\approx$35of them used Cloudflare as a CDN provider. We identified Hostinger’s parking DNS nameservers as the second most prominent DNS provider, as accounting for $\approx$8of the transient domains, followed by a long tail of popular DNS hosting services including NS1, Squarespace, and GoDaddy. These findings suggest that those domains are generally hosted on similar infrastructure of long-lived domains. In contrast, the large presence of parked domains could indicate domains that have already been removed or caught before misuse.
@@ 4.3. Reasons for Early Removals
In private conversations with two top registrars in Table 3, they confirmed that transient domains are mostly likely malicious. The registrars did not comment on exact numbers but anecdotally acknowledged that with few exceptions, reasons for early removal include abuse, account suspensions, or credit card fraud. Some legitimate cases are domain tasting and right of cancellation, but they are exceptionally rare. To investigate indicators of maliciousness of transient domain names, we analyzed how often blocklists manage to flag them. We considered both early-removed newly registered domain names and 42358 transient domains (Table 3). With early-removed, we refer to newly registered domains that were removed before the end of our analysis period. Unlike transient domains, early-removed domains appear in the zone file snapshots but still have a shorter lifespan than the 1-year typical duration of a registered domain due to the likely forced removal by the registrar (Antonia Affinito and Raffaele Sommese and Gautam Akiwate and Stefan Savage and Kimberly Claffy and Geoffrey M. Voelker and Alessio Botta and Mattijs Jonker, 2022). We observed 555491 domains that were deleted before 1 February 2024 (10% of newly registered domains we detected).
@@
Of the 42358 transient domains, at least one blocklist flagged 5% (2123) as malicious. Of these, 5% (105) were flagged on their registration date, 1% (12) were already on the blocklists before their registration date, and 94% (2006) appeared after domain deletion.
@@
Blocklists identify and label as malicious a tiny percentage of domains before their registration dates, which may indicate a re-registration. Additionally, a significant percentage (94%) of transient malicious domains appeared on the blocklists post-deletion, indicating that blocklists do not promptly or in some cases ever detect short-lived malicious domains (Bouwman et al., 2022).
@@ 4.4. Visibility Gap in Domain Detection
We first compared newly registered domains, finding that the SIE NOD feed detected $\approx$5more domains than our method. However, the overlap between the two data sources is $\approx$60%, indicating that each method detects a unique subset of domains. Looking at transient domains identified by the two sources, we find the overlap drops further. A total of 855 transient domains were identified by either one of the two sources. Of these, only 33% were detected by both, with the SIE NOD feed identifying 10% more transient than our method. While the SIE NOD detects a larger number, the disjoint nature of the intersecting set of domains suggests the need to combine them to narrow the visibility gap.
To validate our visibility gap against a source of ground truth, we use the perspective of a mid-sized ccTLD registry (.nl) on transient domains. From November 2023 until January 2024, the registry observed 714 domain names that were deleted in less than 24 hours in their registration system. Of those domains, 334 were registered and deleted such that they were never captured in zone file snapshots. Applying our methodology to this ccTLD, we found only 99 transient domains, or 29.6% of the 334 ccTLD-identified transient domain names. [^5] We were unable to compare the SIE NOD feed to the registry data because we had access only to one day’s worth of data. However, based on our previous results we argue that combining the two data sources would still leave a large visibility gap. This result shows that while we can catch a fraction of these transient domains, researchers still have a huge blindspot regarding intra-day events.
@@ 5. Discussion and Conclusion
In the meantime, transient domains, in which malicious activity tends to dominate, have largely been invisible to the research community. At the same time, research has established that the existing DNS-related threat intelligence community, e.g., blocklist companies, are less effective against new and emerging threats (Bouwman et al., 2022), suggesting the need for new approaches.
@@
Verisign long ago supported rapid zone updates, enabling updates to the .com and .net zone every 5 minutes. The data included domain names, nameservers, IP address additions, deletions and modifications, to “promote security and stability by providing a useful tool to online security companies, ISPs, search engines, financial services companies, and other stakeholders”  (ICANN, 2007), [B]. This service illuminated the activity of bad actors, including phishing, fraud and identity theft (Hao et al., 2013). Concerns about abuse of the data by spammers led to Verisign’s termination of this service. In particular, attackers exploited a brief time window where the zone updates were publicly available, but before domain owners had set up protection of their domain.
@@
Given the inherent advantages that attackers have over defenders, the ineffectiveness of existing uncoordinated countermeasures, and the limited obligations (or ability) of registrars to mitigate harm, [^6] Interisle reports that ICANN’s recently revised contract language does not attribute any duty of care to registrars or registries, and does not address vulnerabilities such as exact-match brand registration and suspicious bulk registration (Dave Piscatello and Colin Strutt, 2023). we see a need to expand transparency and accountability mechanisms.
Registrars and registries who want to establish themselves as serious about security could resurrect the capability to subscribe to rapid zone updates along with a framework to safeguard against abuses, and to learn from history how to mitigate the risk of abuse. While risks of abuse of access to non-public DNS data has created tension and controversy for decades, the CZDS program is testament to the ability to manage these risks. More recently, ICANN has coordinated a model for requesting access to other non-public registration data in a more consistent and standardized format (ICANN, 2023). We advocate for the development of a similar model for data sharing among approved and trusted parties, including operators, law enforcement, and the research community. Such access to rapid zone updates would enable a source of labeled data that would allow application of machine learning techniques not only to security research on systemic harms, but also to anti-abuse efforts of registries, registrars, and law enforcement agencies.
Future Directions
To demonstrate the visibility gap of daily zone snapshots, we proposed a methodology to identify newly registered domains based on CT logs and active measurements, which we released as a public stream (OpenINTEL, 2024). Our results shed light on previously undetected short-lived domains, often related to malicious activities, which we named transient domains. Preliminary analysis and contact with registrars suggest these domains are likely malicious. In the future we plan to expand our measurements beyond DNS infrastructure records, including mail extensions (e.g., SPF, MX, etc.), subdomains, and web-crawling. These measurements, combined with early detection using our methodolgy, can support machine learning-based approaches to proactively identify malicious domains before they do harm. Proactive registries can leverage such use cases to justify and sustain a responsible RZU service.
Acknowledgments
We thank our anonymous shepherd and IMC reviewers for their insightful suggestions and feedback. We thank Cristian Hesselman and Roland van Rijswijk-Deij for their support and insights over time. Additionally, we extend our thanks to Domain Tools and SIDN Labs for providing us with access to passive DNS feeds and registry ground truth. This material is based on research sponsored by the National Science Foundation (NSF) grant OAC-2131987. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of NSF. This research was made possible by OpenINTEL, a joint project of the University of Twente, SURF, SIDN, and NLnet Labs; and operational support from the Twente University Centre for Cybersecurity Research (TUCCR).
@@ Appendix A Ethics
Our ethics review board does not require a specific ethics approval for internet measurements conducted according to community best practices. In measuring the DNS infrastructure of newly registered domains, we followed these best practices by aiming to strike a balance between coverage and slower probing rate. As a result, we decided to scan those domains every 10 minutes, with only three infrastructural query type (A, AAAA, NS), and without enumerating possible sub-labels. Furthermore, we ensured that the address space of our measurement infrastructure had pointers to direct interested users to a page explaining the project’s goals and providing a contact for opting out from the measurements.
Regarding RDAP measurements, we queried approximately 76,000 domains per day. We limited our measurements to a maximum of one domain per second to accommodate possible bursts. This rate is generally acceptable and falls well below most registry rate limits (e.g., 7,200 queries per hour for CentralNic) for RDAP. We distributed measurements across four workers with different IPs, to avoid rare cases of registry RDAP servers enforcing extremely aggressive rate limiting and permanently blocking IPs. To prevent overburdening the infrastructure, we purposefully avoided repeating RDAP measurements in case of failure. We monitored the failure rate to ensure that we did not overload the targeted servers.
@@ Appendix B Appendix: Verisign’s request to provide Rapid Zone updates
The following is a quotation from Verisign’s application to ICANN to provide rapid zone updates as a service  (ICANN, 2007):
@@
  • The service would shed light on the activity of those engaged in domain tasting and expose bad actors. “This [service] elevates bad actors into the light of day.” “All it does is out people who are probably bad guys.”
@@
External counsel and ICANN staff agreed that there are no apparent competition issues that require that this matter be forwarded to relevant competition authorities at this time.